OpenStudy (anonymous):

Which of the following points are NOT on the graph of: y = 2/3x+1 9,7 -3,-1 -6,-3 6,3

OpenStudy (amistre64):

its simplest just to plug the values back into the equation to see if they ring true or false

OpenStudy (amistre64):

for example: y = 2/3x+1 ; take the point (6,3) 3 = 2/3*6+1 3 = 12/3 + 1 3 = 5 is false

OpenStudy (anonymous):

yeah that's correct.none of the choices